OpenStudy (anonymous):

i need to check this gcf of 9 and 16 is 3

OpenStudy (anonymous):

3 is not a factor of 16

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[9=3^2,16=2^4\] so the numbers are "relatively prime" and their greatest common factor is 1

OpenStudy (anonymous):

no wait so the answer is 1

OpenStudy (anonymous):

yes the answer is 1